OpenStudy (anonymous):

A blue bike is $14 less than a red bike. The sum of their prices is $300. How much is the red bike?

OpenStudy (anonymous):

R=B+14 R+B=300 (B+14)+B=300 2B=300-14 B=150-7=143 so your formula would be r=b+14=143+14= can you do the rest?